OpenStudy (anonymous):

A die is rolled. Find each probability. 1. P(less than 3) 2. P(multiple of 3 or 4) 3. P(multiple of 2 or 3)

OpenStudy (anonymous):

2/6 3/6 4/6

OpenStudy (anonymous):

less than 3 {1,2} multiple of 3 or 4 {3,4,6} multiple of 2 or 3 {2,3,4,6}

OpenStudy (anonymous):

For case 1, your looking for the probability of either a 1 or 2 rolling up. There are 6 possible outcomes 1-6. Thus P for 1-2 is equal to \[Probability = Number of outcomes we look for \div Total number of outcomes\] Thus case 1 - 2/6 = 1/3 Case 2 - 3/6 = 1/2 Case 3 - 4/6 = 2/3

OpenStudy (anonymous):

The total number of possibilities is 6. 1. two cases: \[2/6 = 1/3\] 2. three cases (3,4,6) : \[3/6 = 1/2\] 3. four cases (2,3,4,6): \[4/6 = 2/3\]

OpenStudy (anonymous):

Thank all your guy